A pentagon has side angles of 112,93,116,110. Find the missing angle measurement.
First, I added 112+93+116+110 = 431, afterwards I multiplied 431 * 180 = 77,580. Would that be correct?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52787)

.
A pentagon has side angles of 112°,93°,116°,110°. Find the missing angle measurement.
First, I added 112+93+116+110 = 431, afterwards I multiplied 431 * 180 = 77,580. Would that be correct?
---------------------------------------------------------------------

An n-sided polygon has the sum of its interior angles equal to (n-2)*180°.
(See the lesson Sum of interior angles of a polygon in this site).

For example, a triangle has the sum of interior angles equal to 180°.

Particularly, a pentagon has the sum of interior angles equal to (5-2)*180° = 540°.

So, to find the missing angle, you should distract the sum of four given angles from 540°.

See the lesson Solved problems on interior angles of a polygon (Problem 2) in this site.


Answer by MathTherapy(10552)


A pentagon has side angles of 112,93,116,110. Find the missing angle measurement.
First, I added 112+93+116+110 = 431, afterwards I multiplied 431 * 180 = 77,580. Would that be correct?
Why would you multiply the sum of 4 of the 5 interior angles by 180? Who told you to do that? Do you have some sort of resource
(text book, internet, etc.) that will give you some idea of how to approach and determine the answer to this math problem?